Determine whether they system of linear equations has one and only one solution, infinitely many solutions, or no solution. Find all solutions whenever they exist.
2x – 5y = 10
4x – 10y = 20

Answer by stanbon(75887)

Determine whether they system of linear equations has one and only one solution, infinitely many solutions, or no solution. Find all solutions whenever they exist.
2x – 5y = 10
4x – 10y = 20
-------
Notice that both equations have slope = 2/5
The lines are parallel.
Ans: No solutions because the lines do not cross at any point.
===========
Cheers,
Stan H.
==============